### v1

# GVDepth: Zero-Shot Monocular Depth Estimation for Ground Vehicles based on Probabilistic Cue Fusion

Abstract
Generalizing metric monocular depth estimation presents a significant challenge due to its ill-posed nature, while the entanglement between camera parameters and depth amplifies issues further, hindering multi-dataset training and zero-shot accuracy. This challenge is particularly evident in autonomous vehicles and mobile robotics, where data is collected with fixed camera setups, limiting the geometric diversity. Yet, this context also presents an opportunity: the fixed relationship between the camera and the ground plane imposes additional perspective geometry constraints, enabling depth regression via vertical image positions of objects. However, this cue is highly susceptible to overfitting, thus we propose a novel canonical representation that maintains consistency across varied camera setups, effectively disentangling depth from specific parameters and enhancing generalization across datasets. We also propose a novel architecture that adaptively and probabilistically fuses depths estimated via object size and vertical image position cues. A comprehensive evaluation demonstrates the effectiveness of the proposed approach on five autonomous driving datasets, achieving accurate metric depth estimation for varying resolutions, aspect ratios and camera setups. Notably, we achieve comparable accuracy to existing zero-shot methods, despite training on a single dataset with a single-camera setup. Project website: https://gvdepth.github.io/
1 Introduction
Accurate 3D reconstruction is one of the fundamental computer vision challenges, with far-reaching applications across various fields. Unlike multi-view approaches that rely on complex dense matching and pose estimation, learning-based Monocular Depth Estimation (MDE) exploits semantic and geometric cues to infer pixel-wise depth maps from a single image. However, these cues are often specific to a narrow distribution of the training data, resulting in poor generalization to out-of-distribution scenarios.
Figure 1: Proposed Vertical Canonical Representation. We introduce a novel intermediate representation that enables exploitation of underlying planar geometry priors, facilitating learning and generalization across varying camera setups. Under this representation, camera parameters and depth are completely disentangled, enabling depth regression via known depth of the ground plane.
In addition to the environmental domain gap, MDE is particularly sensitive to the domain gap induced by the perspective geometry parameters of the camera system. This issue is present in most supervised [11, 24, 1, 51, 34] and self-supervised methods [55, 52, 14, 15, 54, 53, 42] that are trained on a dataset collected with a single camera system. Models tend to overfit to this perspective geometry bias, leading to significant performance degradation when inferring depth with different cameras. Unfortunately, training a MDE model on data with a wide distribution of perspective geometries is far from trivial. Due to the inherent ambiguity of the MDE problem and entanglement of depth and camera parameters [35], many methods resort to relative depth estimation with unknown shift and scale [8, 43, 36, 48, 49, 21, 46]. Recently, zero-shot metric depth estimation has been achieved with various approaches, including embedding [16], estimating [3, 35] or disentangling the camera parameters [50]. While these methods present impressive results, they require training on data acquired from multiple camera systems, which is not always avaialable or cost-effective.
One specific area with numerous applications across varying fields is MDE for ground-based vehicles and robots. In this context, an additional perspective geometry constraint arises due to the fixed position of the camera relative to the ground plane. This constraint induces an additional depth cue: the vertical image position of the ground-contact point, which has been shown to be a dominant cue in autonomous driving scenarios [10]. The vertical image position cue is closely related to the depth of the ground plane that has recently been leveraged to deliver State-of-the-Art (SotA) depth estimation accuracy in both supervised [47] and self-supervised [42, 4] settings. However, these methods demonstrate limited generalization results, as they are not designed with intention to enable zero-shot estimation.
These findings motivate us to utilize the vertical image position cue and the inherent planarity of the scene in order to enable accurate depth estimation and improve generalization for arbitrary intrinsic and extrinsic camera parameters. Contrary to the object size cue, widely used by the aforementioned zero-shot methods [16, 3, 35, 50], we argue that the vertical image position presents a cue that is more fundamentally rooted in the perspective geometry configuration of the scene. However, without proper modeling this cue is highly susceptible to overfitting to the specific camera system present in the training data [23, 33].
To leverage this cue, while enabling zero-shot generalization, in this paper we propose a method to ensure its consistency for arbitrary camera setups. Specifically, we introduce a novel intermediate depth representation termed Vertical Canonical Representation. Rather than directly regressing depth values, our model predicts vertical image positions of the orthogonal projections of points onto the ground plane, as visualized in Fig. 1. By utilizing the known depth of the ground plane, this approach allows the mapping between the canonical representation and depth space to adapt with varying camera parameters, effectively disentangling the vertical position cue from specific perspective geometry. To further address potential errors in certain regions, we introduce an architecture that adaptively fuses depth predictions from both vertical position and object size cues, guided by dynamically estimated uncertainties.
We conduct comprehensive evaluation of zero-shot MDE in autonomous driving scenarios, thoroughly analyzing the robustness and accuracy of depth estimation with different cues. We demonstrate that our approach is the first to achieve accurate out-of-distribution metric depth estimation when trained solely on a single dataset captured by a single vehicle-camera system. This marks a significant advancement through enabling generalizable metric MDE in data-scarce and specialized scenarios, such as mobile robotics.
2 Related Work
Domain-dependent Monocular Depth Estimation. In a seminal work [11], learning-based MDE has been established as a supervised dense regression task via ground-truth data obtained by reprojection of LiDAR scans. Subsequent works improve upon the original approach by introducing various architectural novelties, such as convolutional models with skip connections between encoder and decoder [24], integration of conditional random fields [25, 51], adversarial training [20, 28] or attention-based architectures [34, 1]. Unfortunately, these methods are specifically fine-tuned for accuracy on well established indoor [39] and outdoor [13] benchmarks, experiencing significant accuracy degradation under distribution shifts [23].
Domain-agnostic Monocular Depth Estimation. The inherent ambiguity and interdependence between camera parameters and depth present challenges for training a robust MDE model that generalizes across a wide range of environments. As datasets are usually collected with a single camera system, models tend to act discriminatively and associate perspective geometry parameters with the pertaining environment. Relative depth estimation methods [8, 43, 36, 48, 49, 46, 21] bypass camera parameter dependency by estimating depth with an unknown scale. While these approaches achieve visually impressive results, their inability to estimate metric depth limits their practicality in real-world applications, often necessitating integration with additional sensors.
ZoeDepth [2] is the first work to show satisfactory metric accuracy across different domains by fine-tuning a pre-trained relative depth estimation model on a single dataset with metric scale. ZeroDepth [16] chooses a different approach by providing the model with embedding of camera parameters, effectively disambiguating the features from the specifics of the camera system. Metric3D [50] addresses the focal length-depth ambiguity by estimating depth in a canonical space defined through focal length denormalization. UniDepth [35] introduces a universal approach, simultaneously predicting both the depth and camera parameters, enabling metric depth estimation for images acquired with unknown sensors. Although these methods demonstrate impressive results, they rely on abundance of training data and computational resources. Furthermore, they are not truly zero-shot, as they tend to overfit to the typically high training resolution, hindering real-time performance on resource-constrained devices that require low-resolution estimation.
Planar Constraints for Depth Estimation. Real-world scenes are often highly regular, featuring numerous planar structures that can serve as additional cues to guide depth estimation [38, 32]. A highly relevant application of MDE is in autonomous driving and mobile robotics, where a fixed relation between the camera system and the ground plane induces additional perspective geometry constraints. It has been shown that models heavily utilize the corresponding ground-contact cues, completely disregarding the object size cue [10], unless models are trained with random crops or more advanced data grafting [33]. Moreover, incorporating ground plane parameterization within the model enhances model accuracy and robustness, both in supervised [47, 23], and self-supervised training [4, 42].
3 Proposed Method
Preliminaries. In this work, we examine an extremely common application domain of MDE: ground vehicles or a robot with a fixed camera configuration. Assuming an ego coordinate system defined at the ground level, we consider a camera system parameterized by intrinsic matrix $\mathbf{K}$ and extrinsic matrix $\mathbf{R|t}$, which transforms points from camera to ego coordinate frame:
$\mathbf{K}=\begin{bmatrix}f_{x}&0&c_{x}\\ 0&f_{y}&c_{y}\\ 0&0&1\end{bmatrix},\mathbf{[R|t]}=\begin{bmatrix}1&0&0&0\\ 0&\cos\theta&-\sin\theta&-h\\ 0&\sin\theta&\cos\theta&0\\ \end{bmatrix},$ (1)
with $\theta,h$ representing camera pitch and height, respectively. We disregard the camera yaw and roll in our parameterization; yaw has no effect on our canonical transformations, while roll is typically negligible in real-world setups.
Problem Statement. We aim to learn a dense mapping $f:(\mathbf{I},\bm{\mathcal{C}})\mapsto\mathbf{D}$, with $\mathbf{I}\in\mathbb{R}^{H\times W\times 3},\mathbf{D}\in\mathbb{R}^{H\times W}$ being RGB image and a correspondending depth map, and $\mathcal{C}$ representing camera parameterization in Eq. 1. Most importantly, the learned mapping should be able to generalize to images acquired with arbitrary camera system. In contrast to existing zero-shot methods [35, 16, 3, 50], which rely on multiple datasets with varying camera parameters, we aim to learn this mapping using data acquired with a single camera system $\bm{\mathcal{C}}_{\text{Train}}$.
3.1 Depth Cues
Figure 2: Depth cues in MDE for ground vehicles. Depth $d$ can be calculated both as a function of imaging size $s$ (with corresponding known object size $S$) and as a function of vertical image position $y$.
Monocular depth estimation relies on semantic and geometric cues in the training data to solve the depth ambiguity of monocular perspective geometry. As illustrated in Fig. 2, one of the possible cues to regress depth $d$ is the relation between object’s real-world size $S$ and imaging size $s$:
$d=f_{y}\frac{S}{s}.$ (2)
However, this cue leads to entanglement between focal length and depth, leading to generalization issues when camera parameters change. Recent zero-shot methods address this issue by embedding [16], estimating [3, 35] or disentangling the focal length [50]. A major drawback of this cue is its reliance on the object’s real world size $S$, which is highly dependent on the particular environmental domain.
For ground vehicles, a fixed position of the camera system relative to the ground introduces an additional perspective geometry constraint – known depth of the ground plane, which is constant for all images acquired by the same camera setup. For each pixel $(u,v)$, depth $d$ is calculated from the intersection of the camera rays $d\,\mathbf{K}^{-1}[u,v,1]^{T}$ and the ground plane $\mathbf{R}^{T}\mathbf{n}-h=0$, with $\mathbf{n}=(0,-1,0)$ being the upward ground normal. This constraint enables depth regression via object’s ground contact point:
$d=\frac{f_{y}h}{\left(H-c_{y}-y\right)\cos(\theta)-f_{y}\sin(\theta)},$ (3)
where $y=H-v$ represents vertical image position of the ground projection, as visualized in Fig. 2. Compared to the object size cue, which is entirely semantic, this cue implicitly encodes the known perspective geometry constraints of the scene, thereby inducing a valuable prior to the ill-posed MDE problem. However, it is highly reliant on accurate and consistent camera calibration.
3.2 Canonical Transforms
Figure 3: Model architecture. Given the input image and corresponding camera parameters, our model predicts two canonical depth representations, $(\mathbf{C_{F}},\mathbf{C_{Y}})$ and corresponding uncertanties $(\mathbf{\Sigma_{F}},\mathbf{\Sigma_{Y}})$. Canonical representations are disentangled from the corresponding camera parameters, enabling separate exploitation of object size cue and vertical image position cue. Canonical representations are then transformed into depth maps $(\mathbf{D_{F}},\mathbf{D_{Y}})$, which are subsequently adaptively fused into a final depth map $\mathbf{D}$ based on estimated uncertanties.
Metric3D [50] proposed a canonical transform disentangling the predicted depth and camera parameters, enabling the use of object size cue for arbitrary camera systems. To achieve this, depth is predicted in canonical space $\mathbf{C_{F}}$ with a canonical focal length $f_{c}$, and post-processed to metric depth $\mathbf{D_{F}}$ by applying Focal Canonical Transform $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\mathbf{D_{F}}=\text{FCT}_{\bm{\mathcal{C}}}(\mathbf{C_{F}})=\frac{f_{y}}{f_{c}}\mathbf{C_{F}}$, and $\bm{\mathcal{C}}=\{f_{y}\}$.
Besides the object size cue, in this paper we propose to leverage the vertical image position cue, effectively encoding the configuration of the scene’s perspective geometry. Unfortunately, this cue also induces entanglement between camera parameters and depth. To enhance cross-dataset generalization, we introduce an additional canonical transform, which enables consistent exploitation of the ground plane prior. Under this formulation, the model estimates a dense map $\mathbf{C_{Y}}$ representing vertical image position of the ground projection point. Metric depth $\mathbf{D_{Y}}$ is then obtained via Vertical Canonical Transform $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$, where $\text{VCT}_{\bm{\mathcal{C}}}(\mathbf{C_{Y}})$ corresponds to mapping $y\mapsto d$ obtainable from Eq. 3. Therefore, this transformation requires knowledge of camera parameters $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$. Crucially, the proposed canonical space remains consistent under variations of camera parameters, facilitating generalization to diverse scenarios.
The proposed VCT is related to the depth of the ground plane, frequently used to enhance MDE accuracy in autonomous driving scenarios [47, 4, 42]. However, these methods exploit the ground-plane constraint exclusively for road pixels, resulting in limited generalization capability.
3.3 Model Architecture
Both cues have their advantages and drawbacks. For example, vertical position cue is highly consistent across environments due to the reduced reliance on semantic information. However, it is highly sensitive to calibration errors, and performs worse as depth increases. An ideal model should leverage both cues, taking into account the confidence for respective image regions. To that end, as visualized in Fig. 3, we design our model to predict depth maps with both cues separately, and then fuse them into a final depth map based on the estimated aleatoric uncertainty.
We use a standard encoder-decoder model with skip connections to process and upsample image features, gaining $\mathbf{F}\in\mathbb{R}^{h\times w\times C}$, where $(h,w)=(\frac{H}{2},\frac{W}{2})$. We forward these features to two very small UNet networks [37], producing our canonical representations $(\mathbf{C_{F}},\mathbf{C_{Y}})$ and accompanying uncertainty estimates $(\mathbf{\Sigma_{F}},\mathbf{\Sigma_{Y}})$. Canonical representations are forwarded to $\text{FCT}(\cdot)$ and $\text{VCT}(\cdot)$, obtaining $(\mathbf{D_{F}},\mathbf{D_{Y}})$, which are then fused into a final depth map $\mathbf{D}\in\mathbb{R}^{h\times w}$:
$\mathbf{D}=\frac{\mathbf{\Sigma_{Y}}\odot\mathbf{D_{F}}+\mathbf{\Sigma_{F}}\odot\mathbf{D_{Y}}}{\mathbf{\Sigma_{Y}}+\mathbf{\Sigma_{F}}}.$ (4)
3.4 Geometric Augmentation
Contrary to zero-shot methods trained on multiple datasets [35, 16, 3, 50], we aim to achieve similar generalization capabilities by training on a single camera system. However, without geometric diversity induced by varying camera parameters, model can simply disregard our proposed canonical transformations. Therefore, we apply geometric augmentations, namely random cropping and resizing, to simulate different camera perspectives. Crucially, compared to [35, 3, 50], we train our model with multiple resolutions, enabling generalization for arbitrary image sizes. A comparable augmentation strategy is suggested in [42, 18], aiming to enhance generalization for arbitrary cameras. However, as we will demonstrate, without the use of canonical transformations, this approach does not effectively transfer across different datasets.
3.5 Optimization
Our main training objective is to guide the fused depth map $\mathbf{D}$ towards the ground-truth depth $\mathbf{D}^{*}$. Given the distance in the $\mathrm{log}$ domain $\mathbf{E}=\log(\mathbf{D})-\log(\mathbf{D}^{*})$, the $\mathrm{SI}_{\mathrm{log}}$ loss [11] is defined as:
$\mathcal{L}_{\mathrm{SI}_{\log}}=\alpha\sqrt{\mathbb{V}[\mathbf{E}]+\lambda\mathbb{E}^{2}[\mathbf{E}]},$ (5)
with $\alpha=10$ and $\lambda=0.15$ as in [1], and $\mathbb{V}$, $\mathbb{E}$ being variance and expectation, respectively.
Additionally, we enhance depth estimation accuracy from both canonical transformations by employing an uncertainty-weighted loss:
$\mathcal{L}_{\text{unc}}=$ $\,\mathbf{\Sigma_{F}}^{-1}\odot||\mathbf{D_{F}-D^{*}}||_{1}+\mathbf{\Sigma_{Y}}^{-1}\odot||\mathbf{D_{Y}-D^{*}}||_{1}$
$+\log(\mathbf{\Sigma_{F}})+\log(\mathbf{\Sigma_{Y}}),$ (6)
which is equivalent to maximizing the log-likelihood, but with L1 norm instead of L2. This approach down-weights regions where each cue exhibits higher uncertainty, encouraging each cue to focus on specific parts of the image where it performs better. To facilitate training stability [22], we predict $\log(\cdot)$ uncertainties, and perform $\exp(\cdot)$ operation accordingly.
Finally, same as in [35], we employ geometric consistency loss. Given two different sets of geometric augmentations of the same image ($\mathcal{T}_{1}$, $\mathcal{T}_{2}$), geometric consistency loss is defined as:
$\mathcal{L}_{\text{con}}=\left\|\mathcal{T}_{2}\circ\mathcal{T}_{1}^{-1}\circ\mathbf{D}_{1}-\operatorname{sg}\left(\mathbf{D}_{2}\right)\right\|_{1},$ (7)
where $\mathbf{D}_{i}$ are estimated depth maps for i’th augmentation, and $\operatorname{sg}(\cdot)$ corresponds to the stop-gradient operation.
Our final loss is a linear combination:
$\mathcal{L}=\mathcal{L}_{\mathrm{SI}_{\log}}+\lambda_{\text{unc}}\mathcal{L}_{\text{unc}}+\lambda_{\text{con}}\mathcal{L}_{\text{con}},$ (8)
with $\lambda_{\text{unc}}=0.5$ and $\lambda_{\text{con}}=0.1$.
4 Experiments
Datasets. We use a diverse set of autonomous driving datasets, each collected with a different camera setup. Specifically, we utilize an aggregation of KITTI [13], DDAD [15], DrivingStereo [45], Waymo [40] and Argoverse Stereo [6] datasets. We split the data into training and validation sets as in the official data splits, utilizing images only from frontal cameras. Due to the missing and sometimes inaccurate information, we calibrate extrinsic parameters, i.e., camera height and pitch, which are required for our canonical representation. Further details are provided in the supplementary materials.
Implementation details. We use a standard encoder-decoder with skip connections [14]. In all experiments we use a ConvNext-Large [27] encoder initialized with ImageNet weights. To provide the model with additional contextual information, we fuse ground plane embedding in the decoder layers [23, 47]. In our depth cue fusion model, secondary U-Nets are designed with a lightweight structure, containing only three convolutional layers for both downsampling and upsampling.
Our models are trained in PyTorch [31] with an AdamW optimizer [29]. Learning rate is progressively reduced with Cosine Annealing from 1e-4 to 1e-5. For all datasets, we use random cropping and resizing to $H\in[200,500]$, with $W$ chosen to preserve the aspect ratio after cropping. To increase resolution diversity in a single batch update, we perform gradient accumulation for 2 training steps. All models are trained for 150000 optimizer steps with an effective batch size of 16, on a single NVIDIA RTX A6000 GPU. Compared to concurrent zero-shot methods [35, 16, 50], this constitutes a fairly simple training setup, both in computational and data resources, demonstrating the efficiency of our approach.
Evaluation details. We use established depth metrics: absolute relative error ($\mathrm{A.Rel}$), root mean squared error ($\mathrm{RMS}$), root mean squared log error ($\mathrm{RMS}_{\log})$, square relative error ($\mathrm{Sq.Rel}$) and accuracy $\mathrm{\delta}_{i}$ under a threshold $1.25^{i}$. Depth is evaluated in the metric range of [0 m, 80 m].
We evaluate GVDepth at an image resolution of $H\times 640$, where $H$ varies according to the dataset to preserve the original aspect ratio. For a fair comparison, domain-specific models are trained and tested using the same image dimensions, which depend on specific dataset. All models are evaluated using a ConvNext-L backbone if available; otherwise, we select the backbone with the most similar complexity. For zero-shot methods, we use the provided checkpoints. For metric zero-shot models which are optimized for a specific training resolution, we report results at both their original training resolution and our testing resolution for a comprehensive evaluation.
4.1 Evaluation of Zero-shot Accuracy
Table 1: Generalization capability comparison. Domain-dependent methods [14, 54, 51, 34] are trained and evaluated on our evaluation setup at an $H\times 640$ resolution. Zero-shot methods [16, 35, 49, 50, 36] are evaluated with provided checkpoints. (†): Shift and scale alignment with ground-truth. (‡): Resizing and padding to native training resolution. Best zero-shot results without resolution alignment are bolded, second best are underlined. In-domain evaluation results are shaded.
Method Train # Train KITTI DDAD
dataset samples $\mathrm{A.Rel}\downarrow$ $\mathrm{RMS}\downarrow$ $\mathrm{RMS}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{\delta}_{2}\uparrow$ $\mathrm{\delta}_{3}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{RMS}\downarrow$ $\mathrm{RMS}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{\delta}_{2}\uparrow$ $\mathrm{\delta}_{3}\uparrow$
Monodepth2  [14] KITTI 23K 9.02 3.94 0.137 91.4 98.3 99.5 37.8 14.6 0.599 20.1 45.7 74.0
DIFFNet  [54] 8.15 3.86 0.129 92.0 98.4 99.6 37.3 14.4 0.563 16.1 42.6 79.1
NeWCRFs  [51] 5.51 2.20 0.084 97.3 99.7 99.9 33.7 15.6 0.571 30.2 62.9 79.2
iDisc  [34] 5.33 2.11 0.081 97.1 99.7 99.9 27.4 13.3 0.461 38.4 74.3 88.4
MiDaS [36] Mix 2M 13.2 5.33 0.199 82.5 95.5 98.5 13.7 8.01 0.319 83.2 95.0 98.1
LeReS  [49] 354K 14.1 4.89 0.190 81.5 95.4 98.3 17.0 7.04 0.331 76.1 93.9 97.6
ZeroDepth  [16] Mix 17M 11.1 3.65 0.154 87.9 96.8 98.9 8.62 5.83 0.156 91.9 97.4 98.8
UniDepth  [35] 3M 13.6 3.13 0.149 88.0 99.3 99.9 13.0 6.62 0.191 84.4 97.1 98.9
Metric3D  [50] 8M 7.91 3.34 0.118 93.0 98.9 99.7 12.0 6.52 0.178 87.7 94.5 98.1
GVDepth KITTI 23K 5.67 2.61 0.090 95.7 99.4 99.9 11.8 8.35 0.251 82.5 93.4 96.5
DDAD 13K 12.3 4.38 0.187 86.0 96.3 98.9 9.67 5.62 0.162 90.3 97.2 98.9
Argoverse 5K 11.2 4.43 0.181 86.1 97.1 99.4 13.5 8.38 0.240 82.6 93.7 96.9
Waymo 156K 10.9 3.23 0.141 89.8 98.4 99.6 13.9 7.68 0.259 86.2 95.2 97.2
DrivingStereo 168K 6.96 3.17 0.110 92.7 98.6 99.7 12.0 7.72 0.215 82.4 93.6 97.6
UniDepth  [35] Mix 3M 5.61 2.34 0.084 97.0 99.6 99.9 15.1 5.82 0.180 89.3 97.7 99.0
Metric3D  [50] 8M 5.41 2.25 0.085 97.2 99.5 99.9 10.6 5.72 0.164 92.4 97.6 98.8
Comparison with State-of-the-Art. We perform extensive evaluation, comparing GVDepth with SotA domain-dependent and domain-agnostic methods. Results for KITTI and DDAD datasets are shown in Tab. 1. Both supervised [34, 51] and self-supervised [14, 54] domain-dependent methods perform well on datasets matching their training domain but overfit to environmental and geometric biases, leading to reduced accuracy in KITTI $\rightarrow$ DDAD transfer. Relative depth estimation models exhibit greater adaptability across domains; however, they require ground-truth scale and shift alignment, limiting their practical applicability. Recent zero-shot metric methods, such as ZeroDepth, UniDepth, and Metric3D, demonstrate strong generalization capabilities but rely on extensive, diverse training datasets, which are not always feasible to obtain. Furthermore, UniDepth and Metric3D suffer accuracy degradation when inferring on resolutions lower than their high-resolution training data. While resizing and padding can address this, these adjustments increase computational demands, limiting their suitability for resource-constrained systems. GVDepth delivers comparable results while being adaptable to arbitrary camera setups, aspect ratios, and image resolutions. Depending on the train-test dataset combination, our method ranks first or second across most metrics, despite training on only a single dataset and using a fraction of the data required by other zero-shot MDE methods. This is particularly impressive, as each dataset features unique environments and camera setups, demonstrating that our model relies on domain-agnostic perspective geometry cues rather than domain-specific semantic cues.
Comparison with ground plane methods.
Table 2: Comparison with methods that leverage known ground plane. Results for KITTI $\rightarrow$ DDAD zero-shot transfer. (†): Results taken from corresponding papers.
Method $\mathrm{A.Rel}\downarrow$ $\mathrm{RMS}\downarrow$ $\mathrm{RMS}_{\log}\downarrow$ $\mathrm{\delta}_{1}\uparrow$
PlaneDepth [42] 32.7 14.0 0.529 37.9
GEDepth [47] 26.1 16.1 - -
Groco  [4] 42.4 15.4 - -
GVDepth 11.8 8.35 0.251 82.5
In Tab. 2 we compare GVDepth zero-shot accuracy with models leveraging the known ground plane for increased performance. GVDepth outperforms competing methods by a significant margin, even though it relies on very similar planar constraint. This highlights the superiority of our Vertical Canonical Representation, which leverages the ground-plane constraint across all image pixels, in contrast to methods like [42, 47, 4] that only improve performance for road pixels.
Qualitative results.
Image Refer to caption Refer to caption
$\mathbf{D}$ Refer to caption Refer to caption
$\mathrm{A.Rel}$ Refer to caption Refer to caption
$\mathbf{\Sigma_{F}}$ Refer to caption Refer to caption
$\mathbf{\Sigma_{Y}}$ Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Qualitative results. Predicted depths, errors and uncertanties for DrivingStereo $\rightarrow$ KITTI evaluation.
In Fig. 4 we illustrate qualitative results of our DrivingStereo $\rightarrow$ KITTI evaluation. Notably, it can be seen that vertical uncertainty $\mathbf{\Sigma_{Y}}$ is lower in closer regions, and decreasing with higher distances. An interesting detail appears in the left image, where the vertical uncertainty adapts to the scene’s leftward slope, indicating that the model is aware of ground perturbations. Additionally, in the right image, the vertical uncertainty is significantly higher for a car lacking a visible ground-contact point, creating ambiguity in its vertical image position cue.
4.2 Ablation studies
Table 3: Ablation of canonical representations. Baseline – standard depth regression with geometric augmentations. Focal – depth regression via Focal Canonical Transform - $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. This is equivalent to Metric3D[50], but trained on our setup. Vertical – depth regression via Vertical Canonical Transform - $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Fusion – depth regression with model as designed in Fig. 3. Training datasets on rows, testing datasets on columns. Best results are bolded, second best are underlined. In-domain evaluation results are shaded. Note: Focal and vertical results presented here are from separate models leveraging $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$, not from submodels in Fig. 3.
Representation KITTI DrivingStereo Waymo Argoverse DDAD
[…]
KITTI Baseline 6.01 2.65 95.2 25.60 8.31 23.0 31.61 10.86 11.0 46.28 15.33 6.0 16.12 8.72 76.3
Focal 5.78 2.78 94.8 12.01 5.77 86.1 19.12 8.30 72.5 10.51 4.86 86.6 14.15 8.41 80.0
Vertical 5.70 2.58 95.5 10.43 5.42 87.3 13.42 7.46 78.6 14.72 7.11 64.7 14.26 9.39 73.6
Fusion 5.67 2.61 95.7 10.24 5.66 87.4 14.34 7.56 78.4 10.45 5.06 84.8 11.81 8.35 82.5
DStereo Baseline 35.62 5.56 20.1 4.82 2.76 97.9 20.21 9.02 65.1 20.62 8.51 50.7 25.32 8.10 47.7
Focal 8.47 3.74 89.8 3.05 1.76 99.5 16.61 8.04 74.2 9.94 4.95 86.8 14.92 7.86 82.0
Vertical 7.42 3.33 92.6 3.07 1.75 99.5 11.82 6.36 85.5 11.41 6.56 84.7 14.71 8.42 80.0
Fusion 6.96 3.17 92.7 3.01 1.76 99.5 12.15 6.50 83.1 9.91 4.92 86.9 12.02 7.72 82.4
Waymo Baseline 18.32 5.65 84.0 22.50 5.98 64.9 4.43 2.86 96.8 28.74 9.51 10.9 24.33 8.92 55.1
Focal 18.21 5.87 80.2 13.81 5.27 85.5 3.62 2.46 98.6 18.41 5.87 85.3 20.41 8.75 66.0
Vertical 8.30 3.17 92.3 11.40 5.25 86.8 3.61 2.48 98.9 12.21 5.12 88.6 11.51 7.34 84.4
Fusion 10.92 3.23 89.8 12.78 5.68 87.1 3.51 2.57 98.8 9.62 4.74 94.1 13.99 7.68 86.2
Argo Baseline 72.31 9.91 7.6 46.32 9.99 20.2 34.41 8.71 34.8 4.72 2.45 97.6 41.92 9.35 25.4
Focal 13.16 6.41 80.2 16.61 10.21 74.6 21.13 9.40 60.6 3.16 1.81 98.6 18.21 10.51 62.6
Vertical 14.60 5.81 84.2 18.22 8.10 75.6 14.98 7.38 72.6 2.95 1.77 98.6 13.86 8.88 81.5
Fusion 11.21 4.43 86.1 12.81 7.81 82.1 12.92 7.38 78.4 3.14 1.82 98.7 13.48 8.38 82.6
DDAD Baseline 59.42 10.17 5.6 18.61 6.72 77.0 26.61 8.40 30.9 31.81 9.98 15.1 10.98 5.86 88.2
Focal 12.45 4.07 86.0 14.31 5.64 83.1 21.91 7.60 53.9 15.01 6.16 83.1 10.91 5.63 88.7
Vertical 22.21 6.17 72.5 17.00 6.45 79.1 15.51 6.27 80.6 14.10 5.96 83.1 10.20 5.73 89.1
Fusion 12.31 4.38 86.0 14.21 6.31 83.8 16.61 7.36 77.6 12.66 5.76 84.9 9.67 5.62 90.3
Evaluation of canonical representations. In Tab. 3 we report extensive evaluation results, spanning different representations and train-test dataset combinations. Moreover, in Fig. 5 we report qualitative results accompanying the table. By carefully examining the results, we can draw the following conclusions: i) The baseline model struggles to transfer effectively to out-of-distribution data. Although baseline models are trained with diverse perspectives through geometric augmentations, they still rely heavily on environmental cues specific to the training domain ; (ii) Vertical image position cues generalize better than object size cues. Models with our novel $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$ representation generally outperform $\text{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models, except in rare cases, potentially stemming from calibration errors. This is particularly evident with models trained on the Waymo dataset, which encounter difficulties when relying solely on object size cues. The probable cause is that Waymo’s cameras are positioned significantly higher, causing a variation of apparent object sizes due to the change of perspective. In contrast, vertical image position cues are more robust, implicitly encoding camera height within the representation; (iii) Probabilistic fusion improves generalization by integrating both cues. Our fusion strategy achieves the most accurate predictions by adaptively weighting the more reliable cue. For example, when object size cues are more accurate (e.g. DDAD $\rightarrow$ KITTI), the model adjusts accordingly, and the same applies for vertical position cues (e.g. Waymo $\rightarrow$ Argoverse).
KITTI Waymo DDAD
Refer to caption Refer to caption Refer to caption Refer to caption
Baseline Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Focal Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Vertical Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Fusion Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
[…]
Figure 6: Error per object class. Vertical – model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Focal – model with $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. Both models are evaluated on DrivingStereo $\rightarrow$ KITTI zero-shot transfer.
Ablation of fusion strategy.
Table 4: Ablation of fusion strategies. Mean – simple mean instead of uncertainty fusion as in Eq. 4. L1 – loss without uncertainty weighting in Eq. 6. All models are trained on DrivingStereo.
Configuration DrivingStereo KITTI Waymo
$\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$ $\mathrm{A.Rel}\downarrow$ $\mathrm{\delta}_{1}\uparrow$
Mean 3.05 99.4 8.21 91.1 14.01 82.8
L1 3.09 99.4 8.02 91.4 13.23 82.9
Fusion 3.07 99.5 6.96 92.7 12.15 83.1
In Tab. 4 we examine the efficacy of different fusion strategies. Zero-shot transfer results demonstrate that our adaptive fusion weighted by aleatoric uncertainties leads to superior generalization performance.
Generalization in specific regions. The strong quantitative generalization results of vertical position cues do not fully capture their performance. For instance, a model with $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ might only perform better on road pixels due to its implicitly encoded planarity. However, as shown in Fig. 6, this is not the case: the $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ model exhibits accurate generalization across a wide range of object classes.
Figure 7: Sensitivity analysis. All models are evaluated on DrivingStereo $\rightarrow$ KITTI. Histograms represent data distribution. a) $\mathrm{A.Rel}$ of $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ and $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$ models depending on ground plane modulation severity. Ground plane modulation is expressed as $\mathrm{A.Rel}$ of ground depth calculated from camera parameters via Eq. 3 and ground-truth depth of the road. b) Ratio of depth errors $(\mathrm{A.Rel}(\mathbf{Z_{Y}})/\mathrm{A.Rel}(\mathbf{Z_{F}})$ and uncertainties $(\mathbf{\Sigma}_{Y}/\mathbf{\Sigma}_{Z})$ depending on depth bin.
Moreover, we examine the effects of ground plane inconsistencies and perturbations on the depth estimation accuracy. A reasonable assumption would be that the $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ model is more sensitive to these perturbations. However, Fig. 7(a) illustrates that error for both models increases in similar pattern as ground perturbation increases. This indicates that our $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ model can adapt and recognize perturbations which are present in the training data, atleast to the similar degree as competing methods.
Finally, in Fig. 7(b) we analyze the errors and uncertainties of both cues across discrete depth bins. As shown, the error ratio skews towards $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ in regions closer to the camera, which is expected. As depth increases, the object size cue becomes more accurate due to the reduced reliability of the vertical position cue at mid-range distances. However, for farther distances, $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ once again achieves lower error. We hypothesize that this is because $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ implicitly encodes the horizon level, aiding depth estimation for distant objects. Additionally, the estimated uncertainties follow a similar pattern, demonstrating the effectiveness of our adaptive fusion module.
5 Conclusion and Future Work
In this work, we present GVDepth, a novel MDE model capable of accurate zero-shot transfer across varying environmental domains and camera setups. Our approach is tailored for autonomous vehicles and mobile robotics, where a fixed relation between the camera system and the ground plane introduces useful perspective geometry constraints. To leverage these constraints for zero-shot generalization, we proposed a Vertical Canonical Representation that uses known depth of the ground plane to enforce invariance to specific perspective geometry, resulting in enhanced learning and generalization. Additionally, we introduced a novel architecture that adaptively fuses depths obtained with object size and vertical image position cues, guided by dynamically estimated uncertanties.
Systematic ablation studies across five autonomous driving datasets demonstrated that our representation significantly improves generalization compared to object size cues, commonly used in existing zero-shot methods. Overall, our approach achieves generalization accuracy comparable to or exceeding state-of-the-art zero-shot models, all while training on a single dataset and requiring significantly fewer computational and data resources. In future work, we plan to extend this approach to multi-dataset and large-scale learning, moving toward a comprehensive foundation model for MDE in autonomous driving and mobile robotics.
Acknowledgment. This research has been funded by the H2020 project AIFORS under Grant Agreement No 952275.
Supplementary Material
A Model architecture
Figure 5: Architecture of our decoder blocks. We adopt a standard upsampling architecture commonly used in U-Net encoder-decoder models with skip connections. The key distinction in our approach lies in incorporating the known depth of the ground plane as positional embeddings, similarly as in [47, 23]. This technique enriches the model with additional contextual information about the camera setup, enhancing its spatial awareness. $\mathcal{F}_{D}^{i}$ denotes the i-th decoder feature map, $\mathcal{F}_{E}^{i}$ the i-th encoder feature map, while $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$ represents the camera parameters. Each convolution operation in the architecture is followed by a GELU activation function [19].
In this work, we employ a fully convolutional encoder-decoder architecture with skip connections. While vision transformer (ViT)-based encoders, such as DINOv2 [30], often achieve superior accuracy, their advantages typically rely on large-scale training. A similar fully convolutional design is adopted in Metric3D [50], demonstrating the effectiveness of such models for generalizable monocular depth estimation (MDE). Given our focus on single-dataset training, we prioritize convolutional backbones to reduce computational complexity while maintaining robust generalization performance.
Decoder architecture details. Our decoder, illustrated in Fig. 5, consists of four blocks that progressively upsample and fuse encoder features from a resolution of $(\frac{H}{32},\frac{W}{32})$ to $(\frac{H}{2},\frac{W}{2})$, while integrating 16-dimensional ground plane embeddings, similarly as in [47, 23]. Decoder channels dimensions are $\{756,512,256,128\}$. Our fusion module processes $(\frac{H}{2},\frac{W}{2})$ resolution feature maps $\mathbf{F}$ with two compact U-Net-like networks [37], each containing 3 downsampling and upsampling blocks, producing two final feature representations. These feature representations are processed with two convolutional blocks which predict our canonical representations $(\mathbf{C_{F}},\mathbf{C_{Y}})$ and accompanying uncertainty estimates $(\mathbf{\Sigma_{F}},\mathbf{\Sigma_{Y}})$. After uncertainty-based fusion, final depth map is bilinearly upsampled to $(H,W)$ resolution.
Figure 6: Vertical Canonical Representation regression. Visualization of regression boundaries of proposed Vertical Canonical Representation. We predict vertical image position of ground plane projection in range $[0,y_{MAX}]$. Image is virtually extended beyond bottom boundary to enable depth estimation for pixels with ground-plane projection below camera field-of-view. For stability reasons, regression is bounded to $y_{MAX}$, which corresponds to $d_{MAX}=80$.
Vertical Canonical Representation details. For convenience, we restate the equation used in Vertical Canonical Transform $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$:
$d=\frac{f_{y}h}{\left(H-c_{y}-y\right)\cos(\theta)-f_{y}\sin(\theta)}.$ (9)
Upon closer inspection, it becomes evident that mapping $y\mapsto d$ can lead to unstable training if not properly regularized. Specifically, at $y$ corresponding to the horizon level of a given camera setup, the depth $d$ approaches infinity. Furthermore, for any $y$ above the horizon, the depth $d$ becomes negative. To address this, as visualized in Fig. 6, we bound the regression of $y$ to $y_{MAX}$, derived from the inverse mapping from $d_{MAX}=80$. Additionally, to accommodate depth regression for objects with ground-plane projections below the camera’s field of view, we virtually extend the image. In practice, this corresponds to increasing $H$ in Eq. 9 accordingly.
B Model Complexity
Table 6: Model architecture details. Encoder architectures number of parameters and execution time for models used in this work. Execution time is measured for resolution 416x640, on a single NVIDIA RTX A6000 GPU.
Method Encoder Params # Execution time
Monodepth2 [14] ResNet-50 [17] 34M 12.70 ms
DIFFNet [54] HRNet-18 [41] 11M 29.42 ms
NeWCRFs [51] Swin-L [26] 270M 62.56 ms
iDisc [34] Swin-L [26] 208M 120.51 ms
PlaneDepth [42] ResNet-50 [17] 39M 15.21 ms
MiDaS [36] ResNeXt-101 [44] 105M 26.51 ms
LeReS [49] ResNet-50 [17] 52M 20.26 ms
ZeroDepth [16] ResNet-18 [17] 232M 238.48 ms
Metric3D [50] ConvNext-L [27] 203M 10.74 ms
UniDepth [35] ConvNext-L [27] 239M 55.34 ms
GVDepth ConvNext-L [27] 228M 21.24 ms
In Tab. 6 we present the architectural details of the models used in this work. While most results align with expectations, there are a few notable outliers worth discussing. First, ZeroDepth exhibits the highest inference times, despite utilizing the least complex backbone among all methods. Essentially, ZeroDepth shifts complexity from the pre-trained backbone to its proprietary self-attention layers in the decoder. This trade-off limits its ability to fully leverage the benefits of large-scale pretraining, which may explain its subpar accuracy compared to UniDepth, Metric3D, and GVDepth, even though it uses significantly more data during training.
On the other hand, Metric3D achieves remarkably low latency in depth prediction despite employing a relatively complex backbone. This efficiency likely stems from predicting depths at a reduced $(\frac{H}{4},\frac{W}{4})$ resolution, which is subsequently upsampled to the original resolution Lastly, GVDepth stands out for its lightweight design and low inference time, even though it uses a relatively complex backbone. Its impressive generalization performance is attributed to our novel methodology rather than reliance on model or data scaling. In future work, we will investigate more complex transformer architectures, which were costly to train on our current computational setup.
Figure 7: Resolution adaptability on KITTI dataset. $\mathrm{A.Rel}(\%)$ of UniDepth [35] and GVDepth as height resolution $H$ decreases from base resolution 462. Computational complexity measured by inference time on NVIDIA RTX A6000 is depicted with different circle radius. GVDepth results are from DrivingStereo $\rightarrow$ KITTI model. Note: GVDepth is adaptable only to resolutions used during training.
Why resolution adaptability matters. In our main experiments, we tackled the issue of resolution adaptability in current state-of-the-art (SotA) zero-shot methods. For instance, UniDepth [35] requires resizing and padding input images to a fixed resolution of 416×640 to achieve optimal accuracy. However, when applied to datasets like KITTI, this approach results in the model processing nearly twice as many pixels as necessary due to padding. This significantly increases computational complexity without adding any meaningful detail to the image.
In contrast, GVDepth is designed to be fully resolution-agnostic, adapting seamlessly to the native resolution of the input image. As shown in Fig. 7, the error rate of GVDepth remains largely unaffected even as the input resolution decreases. This feature is especially advantageous for real-time systems, where dynamically adjusting image resolution is one of the simplest and most effective ways to control computational complexity.
C Dataset details
In this work, we use KITTI [13], DDAD [15], DrivingStereo [45], Waymo [40] and Argoverse Stereo [6] datasets, both for training and evaluation. For KITTI dataset, we evaluate all models on commonly used Eigen split [11] with Garg crop [12], resulting in 23158 training images and 652 testing images. On DDAD dataset we use the official training and validation split, with 12650 and 3950 images, respectively. Since Waymo, DrivingStereo, and Argoverse Stereo are not widely used for MDE evaluation, we simplify the process by creating custom dataset splits. The resulting training splits consist of {156K, 168K, 5K} samples, while the corresponding testing splits contain {5K, 5K, 500} samples, respectively.
D Camera setup calibration
In this section, we provide additional details about our camera calibration procedure. Our proposed Vertical Canonical Transform $\text{VCT}_{\bm{\mathcal{C}}}(\cdot)$, as indicated in Eq. 9, requires the knowledge of camera parameters $\bm{\mathcal{C}}=\{f_{y},c_{y},h,\theta\}$. Here, for all datasets, $f_{y}$ and $c_{y}$ are usually known up to the reasonable error induced by the calibration procedure. However, for certain datasets, camera height $h$ and camera pitch $\theta$ are either unknown, or not properly calibrated. To remain consistent throughout this work, we recalibrate the extrinsic parameters for each dataset, with details provided in Algorithm 1.
Algorithm 1 Estimate Camera Height $h$ and Pitch $\theta$.
1:RGB images $\{I_{i}\}_{i=1}^{N}$, ground-truth depth maps $\{D_{i}\}_{i=1}^{N}$
2:Median camera height $h_{\text{median}}$ and pitch $\theta_{\text{median}}$
3:Initialize empty sets $\mathcal{H}=\emptyset$, $\Theta=\emptyset$
4:for $i=1$ to $N$ do
5:     Run semantic segmentation model on $I_{i}$ to acquire road pixels $\mathcal{P}_{i}$
$\mathcal{P}_{i}=\{(u,v)\mid(u,v)\text{ belongs to the road in }I_{i}\}$
6:     Extract depths $\{D_{i}(u,v)\mid(u,v)\in\mathcal{P}_{i}\}$
7:     Filter road pixels based on depth:
$\mathcal{P}_{i}^{\text{filtered}}=\{(u,v)\in\mathcal{P}_{i}\mid D_{i}(u,v)<20\}$
8:     Project filtered road pixels to 3D points:
$\mathcal{Q}_{i}=\left\{\begin{pmatrix}X\\ Y\\ Z\end{pmatrix}\Bigg{|}\begin{aligned} X&=\frac{(u-c_{x})Z}{f_{x}},\\ Y&=\frac{(v-c_{y})Z}{f_{y}},\\ Z&=D_{i}(u,v)\end{aligned},\forall(u,v)\in\mathcal{P}_{i}^{\text{filtered}}\right\}$
9:     Fit $\mathbf{R}_{i}(\theta_{i})$ (rotation matrix) and $h_{i}$ (camera height) to $\mathcal{Q}_{i}$ with RANSAC plane estimation
10:     Append $h_{i}$ to $\mathcal{H}$ and $\theta_{i}$ to $\Theta$
11:end for
12:Compute median camera height and pitch:
$h_{\text{median}}=\text{median}(\mathcal{H}),\quad\theta_{\text{median}}=\text{median}(\Theta)$
13:return $h_{\text{median}},\theta_{\text{median}}$
For semantic segmentation of the road plane we use the DeepLabv3 model [7].
Table 7: Camera calibration. Camera extrinsics estimated by the calibration procedure described in Algorithm 1.
Dataset $h$[m] $\theta$[°]
Argoverse Stereo [5] 1.678 0.021
DDAD [15] 1.459 -0.519
DrivingStereo [45] 1.739 -0.561
KITTI [13] 1.659 -0.664
Waymo [40] 2.145 -0.331
 Camera height  Camera pitch
Argoverse [5] Refer to caption Refer to caption
DDAD [15] Refer to caption Refer to caption
DrivingStereo [45] Refer to caption Refer to caption
KITTI [13] Refer to caption Refer to caption
Waymo [40] Refer to caption Refer to caption
Figure 8: Calibration histograms. Height and pitch histograms acquired by calibration procedure described in Algorithm 1. Best viewed zoomned in.
Calibration results. Estimated camera height $h$ and camera pitch $\theta$ for each dataset are provided in Tab. 7. Moreover, in Fig. 8 we visualize the histograms obtained with our calibration procedure. Since DrivingStereo [45] and KITTI [13] do not report extrinsic parameters, we use the estimated values in our proposed canonical transform. Furthermore, for the DDAD dataset [15], we observe a significant discrepancy between the official and estimated extrinsics. Therefore, we use our calibrated parameters, as they align more closely with the ground-truth depth. In contrast, for Argoverse Stereo [5] and Waymo [40], the estimated values are consistent with the official calibration.
KITTI calibration issues. We would also like to address the problems with KITTI calibration. As shown in Fig. 8, the pitch histogram for the KITTI dataset does not follow an unimodal distribution. We believe that this highlights inconsistencies in the official KITTI intrinsics calibration across different recording sequences. While this observation is not novel [9], it is rarely, if ever, discussed in the context of MDE. Given that the KITTI dataset remains a cornerstone for MDE evaluation, we believe that this issue should be given increased attention.
Evaluation with per-frame calibration.
Table 8: Evaluation with per-frame extrinsics calibration. Vertical – depth regression via Vertical Canonical Transform - $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Fusion – depth regression with fusion model. Vertical+ and Fusion+ indicate usage of per-frame camera extrinsic calibration during evaluation. Training datasets on rows, testing datasets on columns. Best results are bolded. In-domain evaluation results are shaded.
Representation KITTI DrivingStereo
[…]
KITTI Vertical 5.70 2.58 95.5 10.43 5.42 87.3
Vertical+ 5.72 2.62 95.6 10.45 5.46 87.1
Fusion 5.67 2.61 95.7 10.24 5.66 87.4
Fusion+ 5.61 2.60 95.8 10.22 5.60 87.5
DStereo Vertical 7.52 3.33 92.6 3.07 1.75 99.5
Vertical+ 7.31 3.28 92.7 2.78 1.61 99.6
Fusion 6.96 3.17 92.7 3.01 1.76 99.5
Fusion+ 6.85 3.15 92.7 2.91 1.74 99.6
Our per-frame calibration procedure offers potentially higher accuracy than median-filtered results. This is because it better captures ground plane perturbations and vehicle dynamics, which cause slight variations in camera pitch. In Tab. 8, we present evaluation results using per-frame calibration. While this approach leads to slight performance improvement, the gain is marginal, possibly due to the noise in our calibration process. Although such additional information about vehicle dynamics and perturbation could theoretically be integrated into a real system using localization with visual odometry or sensor fusion, we exclude these results from the main paper since competing methods do not leverage this information.
Figure 9: Calibration sensitivity. Sensitivity to Gaussian noise perturbations in extrinsic parameters evaluated on KITTI dataset. Vertical – depth regression via Vertical Canonical Transform - $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$. Fusion – depth regression with fusion model. Train/test dataset combination for both models is DrivingStereo $\rightarrow$ KITTI.
Calibration sensitivity. Models involving canonical mapping via $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ are inherently sensitive to inaccuracies in camera extrinsic parameters. In Fig. 9, we evaluate the $\mathrm{A.Rel}$ metric for our Vertical and Fusion model configurations under varying levels of Gaussian noise in the extrinsic parameters.
While both models exhibit sensitivity to calibration noise, the Fusion model demonstrates a smaller error increase due to its uncertainty-based fusion with depth from $\mathrm{FCT}_{\bm{\mathcal{C}}}(\cdot)$. However, the Fusion model’s error still increases under noisy conditions, indicating that its uncertainty prediction does not fully compensate for inaccuracies in camera calibration. This is expected, as camera calibration directly affects the canonical mapping performed by $\mathrm{VCT}_{\bm{\mathcal{C}}}(\cdot)$ and is not internally estimated by the model.